# Adaptive Layer Splitting for Wireless LLM Inference in Edge Computing: A Model-Based Reinforcement Learning Approach

### v3

@@ I Introduction
The field of natural language processing (NLP) has recently experienced transformative changes, driven by the rapid advancement of large language models (LLMs) such as GPT-4 [1] and Gemini [2]. These models are highly proficient at generating human-like text [3, 4, 5, 6, 7], catalyzing progress across various domains [8, 9, 10, 11]. Despite the capabilities of LLMs in centralized cloud environments, towards personalized and specialized tasks, LLMs face significant scalability and privacy issues [12, 13], which has driven the exploration of edge computing [14, 15] as a complementary paradigm.
Notably, edge computing processes sensitive information locally rather than traversing through a centralized cloud [16, 17]. Consequently, it minimizes the exposure to potential privacy breaches and unauthorized access. Moreover, edge computing allows for a flexible and distributed architecture, and can accommodate allocated computational resources to specific requirements [18, 19]. Therefore, the integration of edge computing and LLMs empowers LLMs with enhanced personalized and domain-specific generative capabilities [20]. However, LLMs’ substantial computational demands often exceed the processing capacities of communication-limited user equipment (UE) in radio access network [21]. Correspondingly, split learning and inference [22, 23, 24] are proposed to jointly use the computing capability of UE and edge nodes (e.g., base stations [BSs]).
Though most of the existing works [25, 26, 22, 27, 28, 23, 29] focus on careful model splitting to balance the computational and communication costs, splitting different layers of LLMs is quite distinctive, as it has unanimous dimension of intermediate output, leading to the same communication cost. Meanwhile, transmitting tensors between LLM layers across potentially noisy channels could impede LLM inference performance. As validated in this work lately, given different splitting points, the possible loss induced by unreliable wireless channels produces a significantly diverse impact on model performance. Thus, it remains critical to identify the optimal splitting point for model inference, while combating the wireless channel volatility. Such a viewpoint transcends traditional single-step optimization techniques and necessitates a new framework accommodating the sequential nature of decision-making.
@@ II-B Split Inference in Distributed Computing
[38] introduce to split models between clients and servers to avoid transferring raw data, offering new perspectives on distributed computing and privacy-preserving AI. [25, 26, 22] broaden the application of split learning to encompass the distributed deployment of deep neural networks (DNNs) in wireless networks. [27, 28, 23, 29] distribute different portions of DNNs between edge nodes and cloud for collaborative inference task execution, thus reducing response latency and improving the scalability. Different from these existing works, we highlight the impact of different splitting points on inference performance in wireless channels. To our best knowledge, this belongs to the first efforts to address this important issue, and lays the very foundation for further adaptive layer splitting to balance LLM inference performance and UE computation cost.
@@ II-C Reinforcement Learning in Network Optimization
RL emerges as a pivotal tool for optimizing decision-making processes in dynamic and uncertain environments, as highlighted by [44, 45]. [46, 47] demonstrate RL’s efficacy in enhancing performance optimization within distributed networks, highlighting its potential to adapt and respond to evolving environmental conditions. In the context of wireless network environments, [48, 49] leverage RL to address the challenges of resource allocation and network traffic management, showcasing its capability to optimize system performance amidst the fluctuating nature of wireless communications. With its predictive modeling capabilities and remarkable sampling efficiency, MBRL is adept at navigating environments with variable factors [35, 50]. To address computational constraints and wireless channel volatility, our research takes inspiration from MBRL and devises a computation-efficient reward surrogate model to optimize LLM deployment at the edge.
@@ III System Model and Problem Formulation
$h$ Rayleigh fading channel gain
$\sigma_{h}$ Rayleigh fading channel scale parameter
@@ III-A System Model
To characterize the LLM provisioning with model splitting, we primarily consider a system model illustrated in Fig. 2. Without loss of generality, we assume that for an $L$-layer LLM, the first $p$ layers $L_{U}(\cdot)$ are deployed at the UE, and the remaining $L-p$ layers $L_{C}(\cdot)$ are in the edge, where the splitting point $p$ is adjustable. On this basis, for an input $x\in\mathbb{R}^{d_{\text{in}}}$, typically a sequence of tokens in text, a UE transforms $x$ into a higher-dimensional intermediate tensor $y\in\mathbb{R}^{d_{\text{mid}}}$, that is,
@@
where $\theta_{\text{UE}}$ denotes the parameters of layers in $L_{U}(\cdot)$. During this procedure, it inevitably incurs a certain computational load on UEs, which is typically measured in terms of floating-point operations per second (FLOPs). Typically, for the layer $1$ with an input sequence of length $d_{\text{in}}$ with hidden dimension $d_{\text{mid}}$, utilizing a multi-head self-attention mechanism with $\kappa$ heads, the computations for the involved multi-head attention mechanism and feed-forward network components can be obtained as $\text{FLOPs}(L_{1})=\frac{3d_{\text{in}}d_{\text{mid}}^{2}}{\kappa}+\frac{2d_{\text{in}}^{2}d_{\text{mid}}}{\kappa}+d_{\text{in}}d_{\text{mid}}^{2}+8d_{\text{mid}}^{3}$. Thus, the computational load on the UE can be approximated as
$C_{\text{UE}}(p)$ $=\sum_{i=1}^{p}\text{FLOPs}(L_{i})$
$=\frac{3d_{\text{in}}d_{\text{mid}}^{2}}{\kappa}+\frac{2d_{\text{in}}^{2}d_{\text{mid}}}{\kappa}+d_{\text{in}}d_{\text{mid}}^{2}+8d_{\text{mid}}^{3}$
$+(p-1)\cdot\big{(}\frac{5d_{\text{mid}}^{3}}{\kappa}+9d_{\text{mid}}^{3}\big{)}.$ (2)
The intermediate tensor $y$ is transmitted from the UE to the edge over a wireless communication channel. Mathematically, the received signal $\hat{y}\in\mathbb{R}^{d_{\text{mid}}}$ after a Rayleigh fading channel can be represented as:
@@
where $h\sim\text{Rayleigh}(\sigma_{h})$ represents the Rayleigh fading with scale parameter $\sigma_{h}$, and $n\sim\mathcal{N}(0,\sigma^{2})$ represents the noise following a normal distribution with zero mean and variance $\sigma^{2}$. When $h=1$, it degenerates to an additive white Gaussian noise (AWGN) channel. Besides, when the channel gain $h$ falls below a certain threshold $h_{\text{th}}$, the lower signal-to-noise ratio (SNR) and relatively higher bit error rate (BER) implies retransmission, thus exceeding the latency requirement in quality of service (QoS) with a rather high probability. Hence, such a case can be regarded as a packet loss. Accordingly, recalling the formula of Rayleigh distribution, the probability of packet loss can be expressed as
$P(\text{Packet Loss})=P(h<h_{\text{th}})=1-e^{-\frac{h_{\text{th}}^{2}}{2\sigma_{h}^{2}}}.$ (4)
@@ III-B Problem Formulation
Figure 3: Illustrations of the impact on PPL across different layers for various LLMs under (a) high SNR and (b) low SNR in AWGN; (c) low packet loss probability and (d) high packet loss probability under Rayleigh fading.
Beforehand, we investigate the impact of splitting points on the inference performance under different channel conditions and present the corresponding simulation results regarding several mainstream open-source LLMs, including LLaMA2-7B, LLaMA2-13B [5], Mistral-7B [51], Aya-23-8B [52], Openchat-8B [53], and Prem-1B [54], in Fig. 3. Consistent with our intuition, it can be observed from Fig. 3 that for the same settings, a lower SNR or larger packet loss rate generally yields inferior performance (i.e., larger PPL). More interestingly, an earlier model splitting (i.e., a smaller $p$) could worsen the inference performance while the channel conditions significantly affect the performance of a given splitting point. Such an observation is also consistent with the widely recognized fact that earlier layers in LLM are responsible for learning the general or basic features of the training dataset [55]. These findings underscore the importance of strategically selecting the splitting point in the LLM architecture to maintain a desired performance.
On the other hand, high computational load on UEs would lead to increased latency and energy consumption, further undermining the benefits of deploying LLMs in edge environments. As indicated in (2), the computational load on the UE is approximately proportional to the number of layers processed locally. This proportionality yields a contradicting phenomenon that an earlier model splitting worsens the inference performance but ameliorates the computational cost at the computation-limited UE. In other words, in order to minimize the overall system PPL while effectively balancing the computational load on the UE, the problem turns to identifying the optimal splitting point $p$ under volatile channel conditions, that is,
$p^{*}=\arg\min_{p}\left[\text{PPL}(p;\sigma,\sigma_{h})+\lambda\cdot C_{\text{UE}}(p)\right],$ (7)
where $\text{PPL}(p,\sigma,\sigma_{h})$ quantifies the LLM’p inference performance, taking into account the splitting point $p$, the noise intensity $\sigma$, and the Rayleigh fading scale parameter $\sigma_{h}$, which directly affects the packet loss probability. Besides, the weight $\lambda$ balances the trade-off between the inference performance and computational load.
Given the variability of network conditions and the intricacies of real-time decision-making in distributed systems, we reformulate this optimization problem as a sequential decision-making task. In that regard, RL is particularly well-suited for this scenario, due to its capability to adapt to the evolving environment and optimize decisions accordingly.
@@ IV-A The Markov Decision Process
Figure 4: Illustrations of the RL setup, including the LLM, RL agent, and channel noise modules. The RL agent optimizes the splitting point of the LLM by receiving state inputs (noise intensity, Rayleigh fading scale and splitting point), computing action probabilities via the policy network, and updating the policy based on the reward function.
The splitting point adjustment under volatile channels can be formalized as a Markov Decision Process (MDP), consisting of a tuple $\langle\mathcal{S},\mathcal{A},P,R\rangle$. In particular, as illustrated in Fig. 4, the state space $\mathcal{S}$ encompasses the key factors such as the noise intensity $\sigma$, the Rayleigh fading scale $\sigma_{h}$ and the current splitting point $p$, namely, $s=\{\sigma,\sigma_{h},p\}\in\mathcal{S}$. The action space $\mathcal{A}$ includes three possible actions, that is, moving the splitting point upward $+1$ or downward $-1$, or maintaining the current splitting point 00. Mathematically, $a\in\mathcal{A}=\{+1,-1,0\}$. For a time-step $t$, given an action $a_{t}$ under state $s_{t}$, the environment state will transit to the next state $s_{t+1}$ following the transition probability $P$, which is contingent on the selected action and real-time channel conditions. Meanwhile, a reward can be obtained as
@@ IV-C The Reward Surrogate Model for Faster RL
The integration of MBRL into our LLM split optimization scenario significantly enhances the efficiency and effectiveness of our RL approach. Nevertheless, the slow reasoning capability of LLM makes the learning process sluggish. Therefore, inspired by the classical MBRL, which uses a predictive model to simulate the environment, we adopt a surrogate model to approximate the reward function, thereby boosting the learning efficiency. Specifically, we approximate the $\text{PPL}(p;\sigma,\sigma_{h})$ that needs to be computed by running LLM, and computes a DNN-based surrogate model $\widetilde{\text{PPL}}(p;\sigma,\sigma_{h},\theta_{\text{surr}})$ parameterized by $\theta_{\text{surr}}$ to minimize the mean squared error (MSE) as
$\text{MSE}=\mathbb{E}\Big{[}\big{(}\text{PPL}(p;\sigma,\sigma_{h})-\widetilde{\text{PPL}}(p;\sigma,\sigma_{h},\theta_{\text{surr}})\big{)}^{2}\Big{]}.$ (13)
@@
$\qquad\qquad\qquad+\gamma V(s_{t+2+l})-V(s_{t+1+l})\Big{]}.$ (14)
Finally, we summarize the algorithm in Algorithm 1.
@@
32:        Train surrogate model $\widetilde{\text{PPL}}(p;\sigma,\sigma_{h},\theta_{\text{surr}})$ to minimize MSE with (13).

### v4

@@ I Introduction
The field of natural language processing (NLP) has recently experienced transformative changes, driven by the rapid advancement of large language models (LLMs) such as GPT-4 [1] and Gemini [2]. These models are highly proficient at generating human-like text [3, 4, 5, 6, 7], catalyzing progress across various domains [8, 9, 10, 11]. Although LLMs perform well in centralized cloud environments, they face significant scalability and privacy issues [12, 13] in sensitive applications like healthcare and finance [14], which have driven the exploration of edge computing [15, 16] as a complementary paradigm. Notably, edge computing processes sensitive information locally rather than traversing through a centralized cloud [17, 18]. Consequently, it minimizes the exposure to potential privacy breaches and unauthorized access. Moreover, edge computing allows for a flexible and distributed architecture, and can accommodate allocated computational resources to specific requirements [19, 20]. Therefore, the integration of edge computing and LLMs empowers LLMs with enhanced personalized and domain-specific generative capabilities [21]. However, LLMs’ substantial computational demands often exceed the processing capacities of communication-limited user equipment (UE) in radio access network or Internet of Things (IoT) systems [22, 23, 24]. Correspondingly, split learning and inference [25, 26, 27] are proposed to jointly leverage the computing capability of UE and edge nodes (e.g., base stations [BSs]).
Though most of the existing works [28, 29, 25, 30, 31, 26, 32] focus on careful model splitting to balance the computational and communication costs, splitting different layers of LLMs is quite unique, as the intermediate outputs have consistent dimensions, leading to the same communication cost. Additionally, transmitting tensors between LLM layers over potentially noisy channels could hinder LLM inference performance. As validated in this work lately, given different splitting points, the possible loss induced by unreliable wireless channels produces a significantly diverse impact on model performance. These attributes necessitate a shift in focus from optimizing transmission efficiency to managing the computational burden on UE. In LLM deployment scenarios, where the scale and complexity of the models impose significant demands on UE’s limited computational resources, the challenge lies in balancing the computational load without degrading inference performance. Thus, it is critical to identify the optimal splitting point for model inference while combating wireless channel volatility. Such a viewpoint transcends traditional single-step optimization techniques and warrants a new framework accommodating the sequential nature of decision-making.
@@ II-B Split Inference in Distributed Computing
[41] introduce to split models between clients and servers to avoid transferring raw data, offering new perspectives on distributed computing and privacy-preserving AI. [28, 29, 25] broaden the application of split learning to encompass the distributed deployment of deep neural networks (DNNs) in wireless networks. [30, 31, 26, 32] distribute different portions of DNNs between edge nodes and cloud for collaborative inference task execution, thus reducing response latency and improving the scalability.
Different from these existing works [30, 31, 26, 32], LLM splitting presents unique challenges. Due to the large computational demands of LLMs, deploying different layers across UE and edge nodes requires careful consideration of the computational load on the UE. The significant disparity in computational capabilities between heterogeneous devices further complicates the deployment process [47, 48]. Also, splitting at different points within the LLM significantly impacts the overall inference performance, as demonstrated by our experiments in Section 3. To our best knowledge, this belongs to the first efforts to address this important issue, and lays the very foundation for further adaptive layer splitting to balance LLM inference performance and UE computation cost.
@@ II-C Reinforcement Learning in Network Optimization
RL emerges as a pivotal tool for optimizing decision-making processes in dynamic and uncertain environments, as highlighted by [49, 50]. [51, 52] demonstrate RL’s efficacy in enhancing performance optimization within distributed networks, highlighting its potential to adapt and respond to evolving environmental conditions. In the context of wireless network environments, [53, 54] leverage RL to address the challenges of resource allocation and network traffic management, showcasing its capability to optimize system performance amidst the fluctuating nature of wireless communications. With its predictive modeling capabilities and remarkable sampling efficiency, MBRL is adept at navigating environments with variable factors [38, 55]. To address computational constraints and wireless channel volatility, our research draws inspiration from MBRL and develops a computation-efficient reward surrogate model to optimize LLM deployment at the edge.
@@ III System Model and Problem Formulation
$h$ Nakagami-$m$ fading channel gain
$m$ Nakagami-$m$ fading channel shape parameter
$\Omega$ Nakagami-$m$ fading channel spread parameter
@@ III-A System Model
To characterize the LLM provisioning with model splitting, we primarily consider a system model illustrated in Fig. 2. Without loss of generality, we assume that for an $L$-layer LLM, the first $p$ layers $L_{U}(\cdot)$ are deployed at the UE, and the remaining $L-p$ layers $L_{C}(\cdot)$ are in the edge, where the splitting point $p$ is adjustable. Based on this, for an input $x\in\mathbb{R}^{d_{\text{in}}}$, typically a sequence of text tokens, a UE transforms $x$ into a higher-dimensional intermediate tensor $y\in\mathbb{R}^{d_{\text{mid}}}$, that is,
@@
where $\theta_{\text{UE}}$ denotes the parameters of layers in $L_{U}(\cdot)$. This procedure inevitably incurs a certain computational load on UEs, typically measured in floating-point operations per second (FLOPs). Typically, for the layer $1$ with an input sequence of length $d_{\text{in}}$ with hidden dimension $d_{\text{mid}}$, utilizing a multi-head self-attention mechanism with $\kappa$ heads, the computations for the involved multi-head attention mechanism and feed-forward network components can be obtained as $\text{FLOPs}(L_{1})=\frac{3d_{\text{in}}d_{\text{mid}}^{2}}{\kappa}+\frac{2d_{\text{in}}^{2}d_{\text{mid}}}{\kappa}+9d_{\text{in}}d_{\text{mid}}^{2}$. Thus, the computational load on the UE can be approximated as
$C_{\text{UE}}(p)$ $=\sum_{i=1}^{p}\text{FLOPs}(L_{i})$ (2)
$=p\cdot\big{(}\frac{3d_{\text{in}}d_{\text{mid}}^{2}}{\kappa}+\frac{2d_{\text{in}}^{2}d_{\text{mid}}}{\kappa}+9d_{\text{in}}d_{\text{mid}}^{2}\big{)}.$
The intermediate tensor $y$ is transmitted from the UE to the edge over a wireless communication channel. Mathematically, the received signal $\hat{y}\in\mathbb{R}^{d_{\text{mid}}}$ after a Nakagami-$m$ fading channel [56, 57] can be represented as
@@
where $h\sim\text{Nakagami}(m,\Omega)$ represents the Nakagami-$m$ fading with shape parameter $m$ and spread parameter $\Omega$, and $n\sim\mathcal{N}(0,\sigma^{2})$ represents the noise following a normal distribution with zero mean and variance $\sigma^{2}$. In our simulation, each element of the intermediate tensor is independently subjected to a packet loss probability, abstracting each element as a separate packet to capture the impact of noise at a granular level. When $h=1$, it degenerates to an additive white Gaussian noise (AWGN) channel. The probability density function (PDF) of the Nakagami-$m$ distribution for the channel gain $h$ is given by:
$f(h)=\frac{2m^{m}h^{2m-1}}{\Gamma(m)\Omega^{m}}e^{-\frac{mh^{2}}{\Omega}},$ (4)
where $\Gamma(m)$ denotes the Gamma function. The shape parameter $m$ controls the severity of fading. When $m=1$, the Nakagami-$m$ distribution degenerates to a Rayleigh fading channel; at lower values of $m$, the channel experiences more severe fading, leading to greater variability in noise levels. In practical scenarios, particularly when splitting LLMs between UE and base stations, the mobile nature of devices often leads to rapidly changing channel conditions. For instance, as a user moves from an open outdoor area into a building or dense urban environment, the shape parameter $m$ in the Nakagami-$m$ distribution would decrease, reflecting more severe multipath fading and greater noise intensity fluctuations. This dynamic environment necessitates continuous adaptation of the splitting strategy to maintain inference performance under varying noise conditions.
Besides, when the channel gain $h$ falls below a certain threshold $h_{\text{th}}$, the lower signal-to-noise ratio (SNR) and relatively higher bit error rate (BER) imply retransmission, thus exceeding the latency requirement in quality of service (QoS) with a rather high probability. Hence, such a case can be regarded as a packet loss. Accordingly, recalling the formula of Nakagami-$m$ distribution, the probability of packet loss can be expressed as
$P(\text{Packet Loss})$ $=P(h<h_{\text{th}})$
[…]
@@ III-B Problem Formulation
Figure 3: Illustrations of the impact on PPL across different layers for various LLMs under (a) high SNR and (b) low SNR in AWGN; (c) low packet loss probability and (d) high packet loss probability under Nakagami-$m$ fading.
Beforehand, we investigate the impact of splitting points on the inference performance under different channel conditions and present the corresponding simulation results regarding several mainstream open-source LLMs, including LLaMA2-7B, LLaMA2-13B [5], Mistral-7B [58], Aya-23-8B [59], Openchat-8B [60], and Prem-1B [61], in Fig. 3. Consistent with our intuition, it can be observed from Fig. 3 that for the same settings, a lower SNR or larger packet loss rate generally yields inferior performance (i.e., larger PPL). More interestingly, an earlier model splitting (i.e., a smaller $p$) could worsen the inference performance while the channel conditions significantly affect the performance of a given splitting point. Such an observation is also consistent with the widely recognized fact that earlier layers in LLM are responsible for learning the general or basic features of the training dataset [62]. These findings underscore the importance of strategically selecting the splitting point in the LLM architecture to maintain desired performance.
On the other hand, high computational load on UEs would lead to increased latency and energy consumption, further undermining the benefits of deploying LLMs in edge environments. As indicated in (III-A), the computational load on the UE is approximately proportional to the number of layers processed locally. This proportionality yields a contradicting phenomenon that an earlier model splitting worsens the inference performance but ameliorates the computational cost at the computation-limited UE. In other words, in order to minimize the overall system PPL while effectively balancing the computational load on the UE, the problem turns to identifying the optimal splitting point $p$ under volatile channel conditions, that is,
$p^{*}=\arg\min_{p}\left[\text{PPL}(p;\sigma,m)+\lambda\cdot C_{\text{UE}}(p)\right],$ (8)
where $\text{PPL}(p,\sigma,m)$ quantifies the LLM’p inference performance, taking into account the splitting point $p$, the noise intensity $\sigma$, and the Nakagami-$m$ fading shape parameter $m$, which directly affects the packet loss probability. Besides, the weight $\lambda$ balances the trade-off between the inference performance and computational load.
Considering the variability of network conditions and the complexities of real-time decision-making in distributed systems, we reformulate this optimization problem as a sequential decision-making task. In that regard, RL is particularly well-suited for this scenario, due to its capability to adapt to the evolving environment and optimize decisions accordingly.
@@ IV-A The Markov Decision Process
Figure 4: Illustrations of the RL setup, including the LLM, RL agent, and channel noise modules. The RL agent optimizes the splitting point of the LLM by receiving state inputs (noise intensity, Nakagami-$m$ fading shape and splitting point), computing action probabilities via the policy network, and updating the policy based on the reward function.
The splitting point adjustment under volatile channels can be formalized as a Markov Decision Process (MDP), consisting of a tuple $\langle\mathcal{S},\mathcal{A},P,R\rangle$. In particular, as illustrated in Fig. 4, the state space $\mathcal{S}$ encompasses the key factors such as the noise intensity $\sigma$, the Nakagami-$m$ fading shape $m$ and the current splitting point $p$, namely, $s=\{\sigma,m,p\}\in\mathcal{S}$. The action space $\mathcal{A}$ is designed to accommodate a range of possible adjustments to the splitting point, allowing for both fine-grained and more substantial modifications. Specifically, $\mathcal{A}$ includes actions such as moving the splitting point upward or downward by $u$ layers, where $n$ can take values such as $1$, $2$, $3$, and so forth, or maintaining the current position. Mathematically, this can be expressed as $a\in\mathcal{A}=\{-u,\dots,-1,0,+1,\dots,+u\}$. This generalized action space provides the RL agent with the flexibility to optimize the splitting strategy dynamically in response to varying channel environments. For a time-step $t$, given an action $a_{t}$ under state $s_{t}$, the environment state will transit to the next state $s_{t+1}$ following the transition probability $P$, which is contingent on the selected action and real-time channel conditions. Meanwhile, a reward can be obtained as
@@ IV-B Proximal Policy Optimization
During the training process, we employ an experience replay mechanism specifically adapted to the dynamic nature of the channel conditions in our scenario. Notably, state-action pairs, including the current splitting point $p$, noise intensity $\sigma$, and channel fading characteristics $m$, are stored in a replay buffer. At each training step, a mini-batch of these pairs is sampled from the replay buffer to update the policy network, ensuring that the model learns from a diverse set of past experiences under varying network conditions. This technique helps to break the temporal correlations inherent in sequential channel variations, leading to more stable learning.
@@ IV-C The Reward Surrogate Model for Faster RL
The integration of MBRL into our LLM split optimization scenario significantly enhances the efficiency and effectiveness of our RL approach. Nevertheless, the slow reasoning capability of LLM makes the learning process sluggish. Therefore, inspired by the classical MBRL, which uses a predictive model to simulate the environment, we adopt a surrogate model to approximate the reward function, thereby boosting the learning efficiency. Specifically, we approximate the $\text{PPL}(p;\sigma,m)$ that needs to be computed by running LLM, and computes a DNN-based surrogate model $\widetilde{\text{PPL}}(p;\sigma,m,\theta_{\text{surr}})$ parameterized by $\theta_{\text{surr}}$ to minimize the mean squared error (MSE) as
$\text{MSE}=\mathbb{E}\Big{[}\big{(}\text{PPL}(p;\sigma,m)-\widetilde{\text{PPL}}(p;\sigma,m,\theta_{\text{surr}})\big{)}^{2}\Big{]}.$ (14)
@@
Incorporating the surrogate model directly into the reward calculation, the reward at time step $t$ can be redefined as:
$\tilde{R}(s_{t},a_{t})=-\left(\widetilde{\text{PPL}}(p_{t},\sigma_{t},m_{t},\theta_{\text{surr}})+\lambda\cdot C_{\text{UE}}(p_{t})\right),$ (15)
where $\widetilde{\text{PPL}}(p_{t},\sigma_{t},m_{t},\theta_{\text{surr}})$ represents the estimated PPL provided by the surrogate model. This reformulation significantly reduces the training burden by replacing the direct LLM inference with an efficient approximation, enabling more rapid policy evaluation and iteration of an RL policy. From a theoretical standpoint [64], the surrogate model effectively reduces the variance in reward estimation by providing a smoothed approximation of the true reward landscape. This smoothing is particularly advantageous in high-dimensional action spaces, where small perturbations in actions could lead to large fluctuations in PPL if calculated directly.
@@
32:        Train surrogate model $\widetilde{\text{PPL}}(p;\sigma,m,\theta_{\text{surr}})$ to minimize MSE with (14).
@@
$\qquad\qquad\qquad+\gamma V(s_{t+2+l})-V(s_{t+1+l})\Big{]}.$ (16)
Finally, we summarize the algorithm in Algorithm 1.